# Graph Attention Network-based Block Propagation with Optimal AoI and Reputation in Web 3.0

### v1

@@ VIII Conclusion
In this paper, we have focused on enhancing the performance of blockchain-enabled Web 3.0, with a particular emphasis on optimizing block propagation in public blockchains. Specifically, we have introduced the AoI as a data-freshness metric to evaluate the efficiency of block propagation. Then, to ensure reliability during block propagation, we have established the miner reputation mechanism based on the subjective logic model, which can calculate miner reputations and prevent the new block from propagating to miners with lower reputations. Moreover, to achieve block propagation optimization, we have proposed a GAT-based reliable block propagation optimization model to minimize the overall AoI of block propagation with the given reputation constraint, thus obtaining the optimal block propagation trajectory. Finally, numerical results have demonstrated that compared with conventional routing algorithms, the proposed scheme can minimize the overall AoI and ensure that miners have high reputation values during block propagation, contributing to enhanced efficiency and reliability of block propagation in public blockchains.

### v2

@@ VIII Conclusion
In this paper, we have focused on enhancing the performance of blockchain-enabled Web 3.0, with a particular emphasis on optimizing block propagation in public blockchains. Specifically, we have introduced the AoB as a data-freshness metric to evaluate the efficiency of block propagation. Then, to ensure reliability during block propagation, we have established the miner reputation mechanism based on the subjective logic model, which can calculate miner reputations and prevent the new block from propagating to miners with lower reputations. Moreover, to achieve block propagation optimization, we have proposed a GAT-based reliable block propagation optimization model to minimize the overall AoB of block propagation with the given reputation constraint, thus obtaining the optimal block propagation trajectory. Finally, numerical results have demonstrated that compared with conventional routing mechanisms, the proposed scheme can minimize the overall AoB and ensure that miners have high reputation values during block propagation, contributing to enhanced efficiency and reliability of block propagation in public blockchains.